# Gradient-Free Adaptive Global Pruning for Pre-trained Language Models

### v1

@@ 4.3 Algorithm Design
In this work, we consider the alternating method to optimize our Eq. 5, i.e., optimize each variable while keeping the rest fixed. The careful and elaborate design of our Eq. 5 allows us to derive a closed-form solution for each of the variables:
Pruning Weight.
@@
First consider optimizing Eq. 5 with respect to $\mathbf{M}_{\ell}$ and $\widehat{\mathbf{W}}_{\ell}$. For each linear layer $\ell$ in a FFN module, the optimal solution minimizes $\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$. To solve it, the first step is to write $\mathbf{z}_{\ell}$ as $\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$, where $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$ ($\dagger$ denotes the pseudo-inverse.). We get $\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$ by plugging $\mathbf{z}_{\ell}$ back in original loss. This aligns with the pruning objective of Eq. 2 and can be analytically solved by existing pruning solver e.g., SparseGPT. The superscript of “$pre$” for $\boldsymbol{a}_{\ell-1}$ is omitted in this section for simpler notation.
Updating Activation.
@@
Minimization for $\boldsymbol{a}_{\ell}$ is a simple least-squares problem similar to weight pruning. However, in this case, the matrix $\boldsymbol{a}_{\ell-1}$ appears in two penalty terms in Eq. 5, so we must minimize $\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}$ for $\boldsymbol{a}_{\ell}$, holding all other variables fixed. By following a very similar idea to Ridge regression, the new value of $\boldsymbol{a}_{\ell}$ is given by:
$(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\phi_{\ell}(\mathbf{z}_{\ell})),$ (6)
@@
The update for $\mathbf{z}_{\ell}$ requires minimizing the following loss:
@@
$\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (7)
@@
This problem can be challenging because of the potentially non-linear function $\phi_{\ell}(\cdot)$. Fortunately, as the non-linearity $\phi_{\ell}(\cdot)$ between two consecutive FFN layers in the transformer is typically ReLU, which works entry-wise on its argument, the entries in $\mathbf{z}_{\ell}$ are decoupled. As a result, solving Eq. 7 is particularly easy when $\phi_{\ell}$ is piece-wise linear (as in the case of ReLU). It can be solved in closed-form following a simple if-then logic. Specifically, one only needs to compute two solutions as:
@@
$\begin{split}&\mathbf{z}_{\ell}^{(1)}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre},\\ &\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},\end{split}$ (8)
where the first solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are negative (reduced to zero by ReLU), and the second solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are non-negative (i.e., $\phi_{\ell}(\mathbf{z}_{\ell})=\mathbf{z}_{\ell}$). For more sophisticated choices of $\phi_{\ell}$, including smooth sigmoid curves, the problem can be solved quickly with a lookup table of pre-computed solutions because each entry only depends on two inputs.
Remark 4.2 (Convergence to global optimum with AdaGP).
@@
Consider the unconstrained problem given in Eq. 5, we observe that, if the activation function $\phi$ is ReLU, (1) the objective function is convex with respect to each variable when all others are fixed, and (2) closed-form solutions exist for each of the variable updates in the alternating optimization scheme. The proposed algorithm resembles a multi-block alternating direction method of multipliers (ADMM) that is shown to converge to the global optimum in many applications.
4.4 Time Complexity Analysis
@@
The proposed AdaGP is composed of three main steps, with the overall time complexity being the sum of the complexities of the individual steps. Consider one epoch of AdaGP, in the step of weights pruning, the time complexity is dominated by the computation of the pseudo-inverse of matrix $\boldsymbol{a}_{\ell}$ with dimensions $n\times h$ (where $n$ is the number of samples and $h$ is the hidden dimension of the transformer) and thus given by $O(nh^{2})$, assuming $n<h$. The exact pruning step, when employing SparseGPT as the solver, has a complexity of $O(h^{3})$. Combining both sub-steps, weight pruning has a total time complexity of $O(h^{3})$. The second step, updating activation, is dominated by a matrix inversion as shown in Eq. 6. Given that the weight matrix $\mathbf{W}_{\ell}$ is of size $h\times h$, this inversion has a time complexity of $O(h^{3})$. The third step, updating outputs, does not involve matrix inversion and its complexity does not exceed that of the previous steps. Thus, the overall time complexity of the algorithm is bounded by $O(h^{3})$. Consequently, our method exhibits a per-epoch time complexity that is on par with SparseGPT.
@@ 5 Experiments
We implement AdaGP in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. Similar to Frantar and Alistarh (2023); Sun et al. (2023), we sparsify each transformer decoder layer sequentially in order, which significantly reduces memory requirements, and only apply global pruning to FFN modules. For the calibration data, we follow Frantar and Alistarh (2023) and use 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning.
Models, Datasets & Evaluation.
@@
Following Frantar and Alistarh (2023), we primarily work with the OPT model family Zhang et al. (2022) to study scaling behavior at various model sizes, but our method can be easily applied to other LLMs such as Llama Touvron et al. (2023). While our focus lies on the very largest variants, we also show some results on smaller models to provide a broader picture. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023). We consider the test sets of raw-WikiText2 Merity et al. (2016) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023). For additional interpretability, we also provide zero-shot accuracy results following the same setup of Sun et al. (2023), which is based on the popular EleutherAI-eval harness Gao et al. (2023).
OPT-1.3b (WikiText2: 14.62; PTB: 20.29; C4: 16.07)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
6420.80
4828.13
3435.99
9998.71
1.1e4
5347.89
8209.13
1.0e4
4917.02
96.68
133.92
48.08
Wanda
21.56
34.77
25.78
142.20
146.76
142.24
5692.65
4751.69
4501.73
15.63
24.04
18.23
SparseGPT
18.04
28.19
21.45
69.67
93.36
60.83
2596.70
2361.86
1363.08
15.11
23.71
17.88
AdaGP
17.82
27.72
20.99
58.92
85.33
58.36
1350.31
1192.36
655.76
14.97
23.40
17.67
OPT-2.7b (WikiText2: 12.47; PTB: 17.97; C4: 14.32)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
1691.74
1237.08
1415.02
1.0e4
7916.69
6050.07
7.9e5
5.3e5
4.7e5
272.34
308.55
267.70
Wangda
88.61
140.09
90.06
6140.81
4746.96
5678.66
3.0e4
3.5e4
2.4e4
13.66
21.67
16.10
SparseGPT
13.79
21.18
16.18
24.32
37.82
25.92
2662.74
2285.01
1776.08
12.62
19.28
15.12
AdaGP
13.82
21.07
16.14
23.87
37.09
24.90
1200.12
759.11
527.70
12.63
19.31
15.12
OPT-6.7b (WikiText2: 10.86; PTB: 15.77; C4: 12.71)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
7054.21
5437.44
4850.25
7937.49
5971.86
6031.54
2.4e4
2.5e4
2.1e4
64.11
92.23
82.67
Wanda
54.95
129.73
116.67
1493.58
1196.93
996.00
2.1e4
2.0e4
1.8e4
11.86
18.54
14.77
SparseGPT
12.27
18.90
15.28
31.04
51.26
29.42
8871.24
5713.57
3797.20
11.30
16.90
13.51
AdaGP
12.16
18.39
14.93
23.96
39.32
26.97
2095.85
1842.48
953.44
11.07
16.73
13.42
OPT-13b (WikiText2: 10.13; PTB: 14.52; C4: 12.06)
Sparsity
70% 80% 85% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
9037.12
7734.58
5909.47
1.1e4
9140.88
6340.22
1.3e4
1.3e4
9087.50
67.07
110.77
52.61
Wanda
30.94
39.26
33.31
4216.04
2894.77
2450.57
1.1e4
1.1e4
7244.96
10.33
15.35
12.54
SparseGPT
10.89
16.35
13.39
21.42
33.62
21.01
8408.03
6380.30
3416.23
10.20
15.14
12.48
AdaGP
10.96
16.57
13.38
19.07
28.77
19.29
2052.27
1536.51
538.61
10.21
15.15
12.48
OPT-30b (WikiText2: 9.56; PTB: 14.04; C4: 11.45)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
8691.40
4769.89
4732.66
8941.81
5292.98
5092.26
3.8e7
3.0e7
1.4e7
517.71
374.54
991.95
Wanda
7766.61
5547.45
5741.74
8770.33
6020.70
7132.20
6354.15
4296.37
4654.27
10.94
16.64
13.50
SparseGPT
9.58
14.41
11.93
16.49
22.01
17.67
5747.87
5169.50
3555.24
9.59
14.05
11.60
AdaGP
9.56
14.40
11.94
15.61
19.64
16.61
3050.63
2712.39
1758.63
9.57
14.04
11.60
OPT-66b (WikiText2: 9.34; PTB: 13.36; C4: 10.99)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
Wanda
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
SparseGPT
9.45
13.64
11.37
28.27
57.41
26.26
7803.10
6594.88
4433.35
9.38
13.37
11.12
AdaGP
9.37
13.66
11.37
16.45
21.00
17.70
7504.17
5644.65
3683.91
9.36
13.36
11.12
Table 1: Perplexity of pruned OPT models in high sparsity regimes (> 60%); the lower the perplexity, the better.
OPT-6.7b
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
66.12
56.03
50.49
65.27
65.72
30.63
27.60
51.69
70%
SparseGPT
61.74
54.87
48.46
63.85
64.31
29.27
25.40
49.70
AdaGP
60.61
54.51
48.8
62.9
64.14
30.03
26.60
49.66
80%
SparseGPT
55.08
48.38
42.22
59.43
57.79
25.85
21.40
44.31
AdaGP
58.69
51.26
43.78
59.67
58.38
26.88
22.00
45.81
90%
SparseGPT
38.53
53.07
26.00
48.07
26.81
21.67
14.40
32.65
AdaGP
46.48
52.71
26.21
51.70
27.44
19.71
13.40
33.95
2:4
SparseGPT
62.11
54.87
49.58
64.4
65.32
30.63
27.60
50.64
AdaGP
63.82
55.96
49.55
64.56
65.70
30.55
27.60
51.11
OPT-13b
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
65.87
57.76
52.44
66.02
67.82
33.46
28.62
53.14
70%
SparseGPT
63.03
54.87
50.89
65.43
67.47
32.85
26.40
51.56
AdaGP
63.85
55.23
50.73
65.67
66.46
31.83
27.20
51.57
80%
SparseGPT
59.72
52.35
46.82
61.48
62.50
31.23
21.80
47.99
AdaGP
60.89
53.07
46.19
62.12
62.21
30.38
23.00
48.27
90%
SparseGPT
47.49
52.71
33.17
51.54
39.98
21.33
17.80
37.72
AdaGP
53.43
52.71
38.19
52.96
46.68
25.26
17.40
40.95
2:4
SparseGPT
64.71
54.51
51.92
65.98
67.47
32.85
28.20
52.23
AdaGP
65.63
55.23
51.93
65.27
67.8
33.45
28.40
52.53

### v3

@@ 4.3.1 SparseLLM on OPT models
In this work, we consider the alternating method to optimize our Eq. 5, i.e., optimize each variable while keeping the rest fixed. The careful and elaborate design of our Eq. 5 allows us to derive a closed-form solution to every subproblem as shown in the following.
@@
Pruning weight. First consider optimizing Eq. 5 with respect to $\mathbf{M}_{\ell}$ and $\widehat{\mathbf{W}}_{\ell}$. For each linear layer $\ell$ in a FFN module, the optimal solution minimizes $\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$. To solve it, the first step is to decompose $\mathbf{z}_{\ell}$ to $\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$, where $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$ ($\dagger$ denotes the pseudo-inverse.) Plug decomposed $\mathbf{z}_{\ell}$ back in original loss and we get $\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$, which aligns with the pruning objective of Eq. 2 and can be analytically solved by existing pruning solver e.g., SparseGPT. The superscript of “$pre$” for $\boldsymbol{a}_{\ell-1}$ is omitted in this section for simpler notation.
@@
Updating activation. Minimization for $\boldsymbol{a}_{\ell}$ is a simple least-squares problem similar to weight pruning. However, in this case, the matrix $\boldsymbol{a}_{\ell-1}$ appears in two penalty terms in Eq. 5, so we must minimize $\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}$ for $\boldsymbol{a}_{\ell}$, holding all other variables fixed. By following a very similar idea to Ridge regression, the new value of $\boldsymbol{a}_{\ell}$ is given by:
$(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\cdot\text{ReLU}(\mathbf{z}_{\ell})),$ (6)
@@
Updating output. The update for $\mathbf{z}_{\ell}$ requires minimizing the following loss:
@@
$\beta\|\boldsymbol{a}_{\ell}-\text{ReLU}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (7)
@@
This problem is non-convex and non-quadratic (because of the non-linear function ReLU). Fortunately, because the ReLU function works entry-wise on its argument, the entries in $\mathbf{z}_{\ell}$ are de-coupled. Solving Eq. 7 is particularly easy for the case of ReLU, as it can be solved in closed form followed by a simple if-then logic. Specifically, one only needs to compute two solutions of a quadratic equation:
@@
$\begin{split}\mathbf{z}_{\ell}^{(1)}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre},\quad\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},\end{split}$ (8)
where the first solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are negative (reduced to zero by ReLU), and the second solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are non-negative.
4.3.2 SparseLLM on LlaMA models
In this section, we introduce how SparseLLM decomposes global pruning into subproblems and solves them iteratively on LlaMA model families. The model architecture of LlaMA can be found in Figure 2. Overall, SparseLLM operates similarly on both LlaMA and OPT models, with the main difference being that LlaMA includes an additional dense linear layer, known as the gate projection layer, and uses the SiLU activation function instead of ReLU.
Pruning weight. In this part, SparseLLM functions almost identically to its operation on OPTs.
Updating activation $\mathbf{a}_{\ell}$. Similarly, for updating $\boldsymbol{a}_{\ell}$, SparseLLM works nearly the same as on OPT. The minimization for $\boldsymbol{a}_{\ell}$ is a simple least-squares problem, akin to weight pruning. However, in this case, the matrix $\boldsymbol{a}_{\ell-1}$ appears in two penalty terms in Eq. 5, necessitating the minimization of:
$\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2},$ (9)
for $\boldsymbol{a}_{\ell}$, with all other variables held fixed. Following a concept similar to Ridge regression, the updated value of $\boldsymbol{a}_{\ell}$ is:
$\big{(}\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I}\big{)}^{-1}\big{(}\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\cdot\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\big{)},$ (10)
Updating output $\mathbf{z}_{\ell}$. Updating $\mathbf{z}_{\ell}$ is somewhat simpler in LlaMA since the activation function applies over the gate projection layer. The update requires minimizing the loss:
$\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (11)
This problem is quadratic when solving for $\mathbf{z}_{\ell}$ with other variables fixed. Through mathematical manipulations, the analytical solution for $\mathbf{z}_{\ell}$ is found by solving a quadratic equation:
$\mathbf{z}_{\ell}^{*}=\frac{(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}+\text{SiLU}(\mathbf{s}_{\ell})\odot\boldsymbol{a}_{\ell}}{\text{SiLU}(\mathbf{s}_{\ell})^{2}+\mathbf{1}},$ (12)
where the division is element-wise and $\mathbf{1}$ denotes the all-one matrix.
Updating gate projection output $\mathbf{s}_{\ell}$. Updating $\mathbf{s}_{\ell}$ involves minimizing:
$\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2}+\alpha\|\mathbf{s}_{\ell}-(\mathbf{M}_{s}\odot\widehat{\mathbf{W}}_{s})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},$ (13)
where $\mathbf{M}_{s}$ and $\widehat{\mathbf{W}}_{s}$ denote the mask and layer weights for the gate projection layer. This problem is non-convex and non-quadratic due to the non-linear SiLU function. However, since SiLU operates entry-wise, the entries in $\mathbf{s}_{\ell}$ are decoupled. Despite LlaMA lacking a simple closed-form solution as in OPT (which uses ReLU), the problem can still be solved quickly and analytically using a lookup table of pre-computed solutions, since each element in $\mathbf{s}_{\ell}$ depends on only three variables.
Remark 4.2 (Global convergence of SparseLLM).
@@ 4.3.2 SparseLLM on LlaMA models
Consider the objective function given by Eq. 5, under the condition that the activation function $\phi$ is ReLU. Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm is guaranteed to converge to a global optimum.
4.3.3 Pruning of MHAs
SparseLLM also prunes other linear layers besides those in FFNs. By following Eq. 4, for each linear layer out of FFN modules, the pruning objective simplifies to $\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}^{pre}\|_{2}^{2}$, which is equivalent (with some simple math) to that of completely local pruning as shown in Eq. 2. Existing LLM pruning solvers such as SparseGPT and Wanda are applicable here.
4.4 Time complexity analyses
@@
The proposed SparseLLM consists of three main steps, with the overall time complexity being the sum of the complexities of these steps. In the weights pruning step, the complexity is dominated by the pseudo-inverse computation of matrix $\boldsymbol{a}_{\ell}$ (dimensions $n\times h$), which is $O(nh^{2})$. Using SparseGPT as the solver, the exact pruning step has a complexity of $O(h^{3})$. The second step, updating activation, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$. The third step, updating outputs, has a lower complexity. Thus, the overall algorithm complexity is bounded by $O(h^{3})$, making our method’s per-epoch time complexity comparable to SparseGPT.
@@ 5 Experiments
Experiments setup. We implement SparseLLM in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. For calibration data, we follow Frantar and Alistarh (2023) and use 128 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning. We followed existing work Frantar and Alistarh (2023); Sun et al. (2023) and pruned all linear layers (in FFN and MHA) to the target sparsity.
@@
Models, datasets & evaluation. We consider the OPT model family Zhang et al. (2022) and LlaMA-2 model family Touvron et al. (2023) in our experiments. We show results on different sizes of models to provide a broader picture for the performances of SparseLLM. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023). We consider the test sets of raw-WikiText2 Merity et al. (2016) (WT2) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023). For additional interpretability, we also provide zero-shot accuracy results following the same setup of Sun et al. (2023), which is based on the popular EleutherAI-eval harness Gao et al. (2023).
Table 1: Perplexity in high sparsity regimes ($\geq 70\%$); the lower the perplexity, the better.
OPT-1.3b (WikiText2 (WT2): 14.62; PTB: 20.29; C4: 16.07)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 6420.80 4828.13 3435.99 9998.71 1.1e4 5347.89 8209.13 1.0e4 4917.02 - - -
Wanda 21.56 34.77 25.78 142.20 146.76 142.24 5692.65 4751.69 4501.73 - - -
SparseGPT 18.04 28.19 21.45 69.67 93.36 60.83 2596.70 2361.86 1363.08 252.81 238.41 146.21
SparseLLM 17.82 27.72 20.99 58.92 85.33 58.36 1350.31 1192.36 655.76 128.83 144.48 106.01
OPT-2.7b (WikiText2 (WT2): 12.47; PTB: 17.97; C4: 14.32)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 1691.74 1237.08 1415.02 1.0e4 7916.69 6050.07 7.9e5 5.3e5 4.7e5 - - -
Wanda 88.61 140.09 90.06 6140.81 4746.96 5678.66 3.0e4 3.5e4 2.4e4 - - -
SparseGPT 13.79 21.18 16.18 24.32 37.82 25.92 2662.74 2285.01 1776.08 91.02 91.79 64.95
SparseLLM 13.82 21.07 16.14 23.87 37.09 24.90 1200.12 759.11 527.70 56.90 77.14 52.77
OPT-30b (WikiText2 (WT2): 9.56; PTB: 14.04; C4: 11.45)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 8691.40 4769.89 4732.66 8941.81 5292.98 5092.26 3.8e7 3.0e7 1.4e7 - - -
Wanda 7766.61 5547.45 5741.74 8770.33 6020.70 7132.20 6354.15 4296.37 4654.27 - - -
SparseGPT 9.58 14.41 11.93 16.49 22.01 17.67 5747.87 5169.50 3555.24 441.35 464.73 209.44
SparseLLM 9.56 14.40 11.94 15.61 19.64 16.61 3050.63 2712.39 1758.63 51.28 73.61 37.99
OPT-66b (WikiText2 (WT2): 9.34; PTB: 13.36; C4: 10.99)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude OOM OOM OOM OOM OOM OOM OOM OOM OOM - - -
Wanda OOM OOM OOM OOM OOM OOM OOM OOM OOM - - -
SparseGPT 9.45 13.64 11.37 28.27 57.41 26.26 7803.10 6594.88 4433.35 6594.37 6329.59 3799.87
SparseLLM 9.37 13.66 11.37 16.45 21.00 17.70 7504.17 5644.65 3683.91 4641.8 5296.93 1618.43
LlaMA-2 7b (WikiText2 (WT2): 5.47; PTB: 37.91; C4: 7.26)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 1058.00 544.43 889.46 6380.27 NaN 4162.92 9498.91 1.02e4 7539.65 - - -
Wanda 2644.22 4040.95 1630.09 1814.01 3376.35 1124.26 5206.93 4607.30 2780.45 - - -
SparseGPT 15.98 302.15 18.58 53.20 803.02 52.57 344.97 2503.82 279.77 68.28 784.79 60.45
SparseLLM 16.15 274.35 18.23 49.96 664.39 47.39 225.23 2233.52 181.56 64.17 667.27 54.56
LlaMA-2 13b (WikiText2 (WT2): 4.88; PTB: 50.94; C4: 6.73)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 30.34 2317.39 28.48 4133.98 4706.65 4112.69 5580.71 5514.22 5090.63 - - -
Wanda 23.42 502.53 32.65 295.29 2340.13 261.15 3003.49 3804.69 1738.73 - - -
SparseGPT 12.98 267.63 15.95 45.59 550.59 45.20 825.99 1410.46 673.33 63.48 660.70 56.29
SparseLLM 12.95 277.76 15.77 36.36 578.35 38.63 646.15 1078.94 466.98 53.71 632.11 50.40